OpenStudy (lacefacexox):

Just wanted someone to check my work since I've been having a hard time with math. Find the point on the terminal side of θ = -3π/4 that has an x coordinate of -1. θ = -3π/4 Find the y-coordinate to find the point. Tanθ = y/x Tan(-3π/4) = y/-1 -3π/4 = ((-sqrt(2)/2),(sqrt(2)/2)) = -1 -1 = y/-1 (-1/1)*-1 = (y/-1)*(-1/1) 1=y So the answer is (-1,1)?

OpenStudy (confluxepic):

It's NEGATIVE 135 degrees.

OpenStudy (confluxepic):

So yeah. Just flip it across the x-axis.

OpenStudy (confluxepic):

@lacefacexox

OpenStudy (lacefacexox):

So the coordinate is (-1/1)?

OpenStudy (confluxepic):

Nope. That's for positive 135 degrees. -3pi/4 = (-)3pi/4 = -135 degrees. Both coordinates will be negative since -135 degrees is in the third quadrant.

OpenStudy (lacefacexox):

(-1,1)***

OpenStudy (lacefacexox):

Oh okay.

OpenStudy (confluxepic):

\(\color{blue}{\text{Originally Posted by}}\) @confluxepic @~ Both coordinates will be negative since -135 degrees is in the third quadrant. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (confluxepic):

@lacefacexox Do you know the unit circle?

OpenStudy (lacefacexox):

I know it, but to be honest, I'm no good at it. Guess its a blessing I'm not a math major. lol

OpenStudy (confluxepic):

Do you see 3pi/4?

OpenStudy (lacefacexox):

Yes, in the second quadrant.

OpenStudy (confluxepic):

Do you see negative 3pi/4? Instead of going counterclockwise, try to go clockwise.

OpenStudy (confluxepic):

http://prntscr.com/c71w1p That should make it easier to see the negative angle.

OpenStudy (lacefacexox):

Ok yes I see it

OpenStudy (confluxepic):

The coordinates are the same, right?

OpenStudy (lacefacexox):

Yeah, except they're both negative

OpenStudy (confluxepic):

Good. So both the x-coordinate and the y-coordinate have to be the same. What is the x-coordinate that they gave you in the equation?

OpenStudy (lacefacexox):

So would this be correct θ= -3pi/4 = -135 degrees Find the y-coordinate to find the point. Tanθ = y/x Tan(-3pi/4) = y/-1 -3pi/4 = ((-sqrt(2)/2),(-sqrt(2)/2)) = 1 1 = y/-1 (-1/1)*1 = (y/-1)*(1/1) -1=y So the answer is (-1,-1)?

OpenStudy (lacefacexox):

x-coordinate was -1

OpenStudy (confluxepic):

CORRECT! I'm proud of you.

OpenStudy (lacefacexox):

yayy. It confused me because I thought with it being -135, I had to use 5pi/4